A water tank has an inlet pipe and a drain
pipe. A full tank can be emptied in 30 minutes if the drain
is opened and an empty tank can be filled in 45 minutes
with the inlet pipe opened. If both pipes are accidentally
opened when the tank is full, then how long will it take to
empty the tank?

Answer by mananth(16946)

A water tank has an inlet pipe and a drain
pipe. A full tank can be emptied in 30 minutes if the drain
is opened and an empty tank can be filled in 45 minutes
with the inlet pipe opened. If both pipes are accidentally
opened when the tank is full, then how long will it take to
empty the tank?
filling ----- 30 minutes
it can fill 1/30 of the tank in 1 minute
...
emptying ---------- 45 mins
it can empty 1/45 of the tank in 1 minute.
...
if both are open effective filling
= 1/30 -1/45
=(3-2)/90
=1/90 of the tank in 1 minute
so it will take 90 minutes to fill the tank when both are open
...
m.ananth@hotmail.ca